How long does it take light to travel through a 3.2-mm-thick piece of window glass? Express your answer in seconds.
Answer #1

time = distance / speed

speed of light in material medium v= c/n

time t = d*n/v and refractive index of glass n = 1.5

time t = {(3.2 mm)*1.5}/(3*108 m/s)= (3.2*10-3 * 1.5 )/ (3*108 ) = 1.6*10-11 seconds

  time t = 1.6*10-11 seconds